Find the general solutions of the systems whose augmented matrices is given?

[ 3 -4 2 0 ]
[ -9 12 -6 0 ]
[ -6 8 -4 0 ]
1
Expert's answer
2020-12-01T03:03:00-0500

The matrix


"\left[ 
\begin{array}{cccc} 
3 & -4 & 2 & 0\\
-9 & 12 & -6 & 0\\
-6 & 8 & -4 & 0
\end{array}
\right]"


is equivalent to the matrix


"\left[ 
\begin{array}{cccc} 
3 & -4 & 2 & 0\\
3 & -4 & 2 & 0\\
3 & -4 & 2 & 0
\end{array}
\right]" after dividing each element of the second row by -3 and each element of the


third row by -2. The last matrix is equivalent to the matrix


"\left[ 
\begin{array}{cccc} 
3 & -4 & 2 & 0\\
0 & 0 & 0 & 0\\
0 & 0 & 0 & 0
\end{array}
\right]" after subtraction from second and third rows the first row.


Therefore, the system is equivalent to the equation "3x_1-4x_2+2x_3=0" which has the following infinitely many solutions:


"\begin{cases}
x_1=\frac{4}{3}x_2-\frac{2}{3}x_3\\
x_2\in\mathbb R\\
x_3\in \mathbb R 
\end{cases}"